OpenStudy (anonymous):

Find the exact value. 1/2cos pi/12+sqrt(3)/2 sin pi/12

OpenStudy (anonymous):

looks like the addition angle formula for sine

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\cos \frac{ \pi }{ 12 }+\frac{ \sqrt{3} }{ 2 }\sin \frac{ \pi }{ 12 }\]

OpenStudy (anonymous):

\[\sin(\frac{\pi}{3})\cos(\frac{\pi}{12})+\cos(\frac{\pi}{3})\sin(\frac{\pi}{12})\]

OpenStudy (anonymous):

no that was wrong sorry

OpenStudy (anonymous):

\[\sin(\frac{\pi}{6})\cos(\frac{\pi}{12})+\cos(\frac{\pi}{6})\sin(\frac{\pi}{12})\]

OpenStudy (anonymous):

that looks better i think

OpenStudy (anonymous):

addition angle formula for sine

OpenStudy (anonymous):

Awesome thanks.